OpenStudy (anonymous):

How can I get the frequency, wavelength, and index of refraction out of the following wave: E=E0*cos[pi*10^15(t-x/.65c)]

OpenStudy (anonymous):

\[E = Eo \cos ( \pi \times 10^{15} (t - x/.65c))\] compare this with the general equation \[E = Eo \cos ( \omega(t - x/v))\] where w is the angular frequency v is the wave velocity... now.. can you proceed?